A plane flies 460 mph in calm air. Flying with the wind, the plane can travel 1560 miles in the same amount of time that it takes to travel 1200 miles flying against the wind. Find the rate of wind.
Answer by macston(5194)

.
w=rate of wind
Time=distance/rate Since times are the same:
1560mi/460mph+w=1200mi/460mph-w Cross multiply.
1560mi(460mph-w)=1200mi(460mph+w)
717600mi^2/h-(1560w)mi=552000mi^2/h+(1200w)mi Subtract 552000mi^2/h from each side.
165600mi^2/h-(1560w)mi=(1200w)mi Add (1560w)mi to each side.
165600mi^2/h=(2760w)mi Divide each side by 2760mi
60 mi/h=w
ANSWER: The rate of the wind was 60 miles per hour.
.
CHECK:
1560mi/460mph+w=1200mi/460mph-w
1560mi/460mph+60mph=1200mi/460mph-60mph
1560mi/520mph=1200mi/400mph
3hrs=3hrs